There are 3 times as many boys as girls. If there are 24 more boys than girls, how many children are there altogether?
Answer by ankor@dixie-net.com(22740)

There are 3 times as many boys as girls.
b = 3g
:
If there are 24 more boys than girls,
b = g + 24
how many children are there altogether?
:
Replace b with 3g in above equation
3g = g + 24
3g - g = 24
2g = 24
g = 12 girls
;
b = 3g
b = 3(12)
b = 36 boys
:
12 + 36 = 48 kids altogether